OpenStudy (anonymous):

describe the steps to simplify the problem \[\left(\begin{matrix}3 \\ x\end{matrix}\right)-\left(\begin{matrix}3 \\ x+1\end{matrix}\right)-\left(\begin{matrix}3 \\ x+2\end{matrix}\right)\]

OpenStudy (saifoo.khan):

\[\left(\begin{matrix}3-3-3 \\ x-x-1-x-2\end{matrix}\right)\]

OpenStudy (saifoo.khan):

\[\left(\begin{matrix}-3 \\ -x-3\end{matrix}\right)\]

OpenStudy (anonymous):

can you write it out, i dont understand.